a motorist travels 160 km hour and 160 km at 100 km/h. what is the average speed of the motorist for the trip?


1
Expert's answer
2020-09-09T10:21:09-0400

Average speed is "v = \frac{L}{t}", where "L" is the total distance covered, and "t" is the time it took to cover it.

During the first part, the distance covered is "L_1 = 160km", and the time is "t_1 = 1 h".

During the second part, the distance covered is "L_2 = 160 km", and the time is "t_2 = \frac{L_2}{v_2}", where "v_2 = 100 \frac{km}{h}". Hence, average speed is:

"v = \frac{L_1 + L_2}{t_1 + t_2} = \frac{L_1+L_2}{t_1 + \frac{L_2}{v_2}} \approx 123.1 \frac{km}{h}".